# Present address: ]Lehrstuhl für Theoretische Chemie, Ruhr-Universität Bochum, 44780 Bochum, Germany

### v3

@@ Abstract
How life started on Earth is an unsolved mystery. There are various hypotheses for the location ranging from outer space to the seafloor, subseafloor or potentially deeper. Here, we applied extensive ab initio molecular dynamics (AIMD) simulations to study chemical reactions between NH3, H2O, H2, and CO at pressures (P) and temperatures (T) approximating the conditions of Earth’s upper mantle (i.e. 10 - 13 GPa, 1000 - 1400 K). Contrary to the previous assumptions that larger organic molecules might readily disintegrate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed without any catalysts and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, urea, and uracil-like molecules. Particularly, our free energy calculations showed that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Moreover, while the pyranose (six-membered-ring) form of ribose is more stable than the furanose (five-membered-ring) form at ambient conditions, we observed the predominant formation of the five-membered-ring form of ribose at extreme conditions, which is consistent with the exclusive incorporation of $\beta$-D-ribofuranose in RNA. We have uncovered a previously unexplored pathway through which the building blocks of biomolecules may have originated in early Earth and other planets. Our findings contribute to an evolving understanding of the fundamental conditions necessary for life to arise.
@@ I Introduction
In those hypotheses, prebiotic chemistry typically happens at non-ambient conditions. The pressure (P) of hydrothermal vents, resulting from the interaction between seafloor and seawater producing hot fluids, is tens of megapascals depending on the depth of water, and the temperature (T) can reach 700 K. Those deep sea vents were considered to have suitable conditions for the origin of life martin2008hydrothermal , but previous studies also suggested that the high temperatures in the vents could readily degrade crucial biomolecules in aqueous solutions miller1988submarine . Shock recovery experiments successfully synthesized various biomolecules, where the pressure can reach tens of gigapascals and the temperature can rise to thousands of kelvins before releasing back to ambient conditions Sekine2024 ; however, due to the dynamic nature of shock waves, samples do not reach thermodynamic equilibrium states, so the shock experiments are usually conducted to mimic the conditions of meteorite impacts on Earth furukawa2009biomolecule . It was commonly assumed that large biomolecules could not persist at both high hydrostatic pressures and high temperatures as found in Earth’s upper mantle, where pressures can reach up to $\sim$13 GPa, accompanied by high temperatures of $\sim$1700 K. In recent years, Otake et al. studied the stability of amino acids and their oligomerization at 1.0–5.5 GPa and 453–673 K, and found that high-pressure conditions inhibited the decomposition of amino acids and small amino acids could be oligomerized up to pentamers; however, their experiments were conducted at the water-poor condition otake2011stability . In aqueous solutions, there seems a strong thermodynamic drive towards the formation of amino acids dimers above 1.5 GPa and temperature above 423 K robinson2021quantifying , in agreement with some experimental observations pedreira2019spontaneous . Previously, the fluids in Earth’s mantle were commonly modeled as simple mixtures of small unreactive volatile molecules, including H2O, CO2, CO, CH4, and H2 zhang2009model . However, recent experimental and theoretical studies have revealed the significant roles played by chemical speciation, aqueous ions, and complexes in the supercritical geofluids within Earth’s lithosphere manning2013chemistry ; pan2013dielectric ; sverjensky2020changing ; pan2016fate ; abramson2017water ; stolte2019large ; stolte2021water . These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules in Earth’s or other planets’ subsurface or even deeper.
@@
It is very challenging to study the origin of life in laboratory settings due to the extreme environments and subsequent analyses. Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular-level insights perez2020prebiotic . Saitta et al. applied ab initio molecule dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller . Goldman et al. applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic . Note that the shock compression simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions Ross2013comment ; Goldman2013reply . AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide . Despite the numerous previous studies, the exploration of prebiotic chemistry at both high hydrostatic pressures and high temperatures largely remains an uncharted territory.
@@
Here, we applied extensive AIMD simulations to study the C-H-O-N fluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found in the transition zone of Earth’s upper mantle. We discovered that a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including formamide, glycine, ribose, urea, and uracil-like molecules. Our free energy calculation suggested that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Our study provides valuable insights into a new potential avenue for the formation of building blocks of life and potentially for the origin of life within Earth’s and other planets’ interior, expanding our understanding of the conditions necessary for life to emerge.
@@ II Results and Discussion
The simulation box initially contained 15 molecules of each of the following species: H2O, H2, CO, and NH3 (Figure S1). Those molecules are typical volatile compounds found in C-H-O-N fluids in reducing environments of Earth’s deep interior zhang2009model ; manning2013chemistry , and are also abundant in interstellar space guelin2022organic . We conducted AIMD simulations at three P-T conditions, 10 GPa and 1000 K, 10 GPa and 1400 K, and 13 GPa and 1400 K, to evaluate how chemical products vary with P and T. Those P-T conditions are typically found at the bottom of Earth’s upper mantle. After more than 400 ps simulations at each P-T condition, we found about 100 different organic species. Here, we mainly focus on the formation of biotically relevant molecules, so we classify the organic molecules into six categories: “CN”, “C2N”, “CN2”, “C2N2”, “C3Nx”, and “C4Nx”, based on the numbers of carbon and nitrogen atoms, as shown in Table 1. Fig. 1 shows the chemical structures of the representative species. Species with more than five carbon atoms were observed; however they are excluded at this stage due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds. The percents of these CN-containing species as a function of time at three different P-T conditions are shown in Figure S3, which suggest that the concentrations of these species gradually stabilize over time.
@@
Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; saladino2015meteorite ; rios2015impact ; ferus2015high ; enchev2021chemical . Thus, we first investigated the formation of formamide molecule. Fig. 3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Figure S4. At 10 GPa and 1000 K, the formation reaction begins with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution. The encounter between an H+ ion and a CO molecule leads to the formation of a C-H bond and an aldehyde group (-CH=O). Subsequently, the aldehyde group reacts with a NH3 molecule to generate a C-N bond, and then, one N-H bond breaks, ultimately producing formamide. At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K. One N-H bond in NH3 first breaks, leading to a dissociative amino group (-NH2). The amino group then reacts with a CO molecule, forming the C-N bond. Later, a water molecule dissociates, providing a proton that ultimately contributes to the formation of the formamide molecule. At 13 GPa and 1400 K, the C-N bond is initially formed through the reaction between NH3 and CO molecules. After a very short time, a dissociated proton from a water molecule is also bonded to the carbon atom, forming a C-H bond. Finally, by releasing a proton from the NH3 group, the formamide molecule is generated. Despite the different formation pathways at three P-T conditions, we observed two common features: first, the reactants always have CO, H2O, and NH3, and second, the reactions must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
@@
For the generation of biomolecules in living matter, the formation of C-N bonds is the key step. We further applied the advanced sampling method (see details in the Methods section) to calculate the free energy profile of the C-N bond formation at three P-T conditions in Fig. 3(b). The collective variable (CV) is the distance between the carbon and nitrogen atoms. Our calculations show that the C-N bonds at 10 GPa, 1000 K and 13 GPa, 1400 K, whose lengths are 1.33 Å and 1.31 Å  respectively, are at the metastable states. The free energy decreases upon dissociation of the C-N bond, indicating that breaking C-N bonds at these two P-T conditions are more thermodynamically favored. The free energy profiles suggest that the energy barriers of the C-N bond dissociation are 17.57 kcal/mol and 16.40 kcal/mol, at 10 GPa, 1000 K and 13 GPa, 1400 K, respectively. As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 59.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition. The free energy result agrees with the data in Fig. 2, where there are more CN-containing species at 10 GPa and 1400 K than at the other two P-T conditions. Additionally, compared with the free energy profile at 10 GPa and 1000 K, there are two more energy minima at the C-N distance of 2.37 Å  and 2.45 Å at 10 GPa, 1400 K and 13 GPa, 1400 K, respectively. The free energy increases upon the dissociation of the C-N pair, indicating that these two minima are thermodynamically stable. The C-N distances are larger than that of a typical C-N covalent bond, which implies the presence of an additional atom connecting the two. This is consistent with a wider variety of organic species at these two P-T conditions. Overall, our free energy calculations suggest that the C-N bond formation is favored at 10 GPa and 1400 K.
@@
After studying the basic chemical products at three P-T conditions, we now turn to investigate the formation of large biomolecules as potential building blocks of life. Previous simulations have extensively studied glycine and the smallest amino acid under various conditions saitta2014miller ; goldman2010synthesis ; wang2014discovering . In our simulations at 10 GPa and 1400 K, we discovered $\alpha$-hydroxyglycine, very similar to glycine. Fig. 4 (a) illustrates the process of forming $\alpha$-hydroxyglycine from formamide. A proton attacked the formamide molecule, leading to the creation of a protonated formamide molecule, also known as an aminohydroxymethyl radical. Then, a CO2 molecule reacts with the protonated formamide, resulting in the “-N-C-C-” backbone through forming a C-C bond. Finally, another proton is captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule. In comparison to glycine, $\alpha$-hydroxyglycine differs by having one hydrogen atom and an “-OH” group instead of two H atoms connecting to $\alpha$-C. Glycine can be obtained if the “-OH” group is replaced by a hydrogen atom. During this process, the formation of the C-H bond is a crucial step. To explore appropriate P-T conditions for the possible synthesis of glycine, we calculated the free energy as a function of the C-H distance in Fig. 4 (b). Considering that the CN-backbone of the “C2N” species at 13 GPa - 1400 K is “-C-N-C-”, different from the CN-backbone in the glycine molecule, we did not consider this P-T condition. We can see that the free energy minima is at 1.07 Å for both 10 GPa, 1000 K and 10 GPa, 1400 K, and the energy barriers to break the C-H bond are 21.91 kcal/mol and 22.04 kcal/mol, respectively, which are very similar. After the C-H bond dissociation, the free energy increases more at 10 GPa and 1400 K (6.76 kcal/mol) than at 10 GPa and 1000 K (2.63 kcal/mol), indicating that the C-H bond is more favored at the former P-T condition.
@@
To study the reaction from $\alpha$-hydroxyglycine to glycine at extreme P-T conditions, we set up a new simulation box containing 5 $\alpha$-hydroxyglycine, 10 H2, and 10 H2O molecules, as shown in Fig. 4(c) and performed AIMD simulations at 10 GPa and 1400 K. Initially, two $\alpha$-hydroxyglycine molecules combine to form one molecule, but later dissociate to dehydroglycine and glyoxylic acid after $\sim$28 ps. As the simulation goes on, the C=N double bond in the dehydroglycine molecule receives two protons and finally dehydroglycine becomes glycine at $\sim$ 38 ps.
@@
We first considered ribose (C5H10O5). The formose reaction is a well-known prebiotic pathway for synthesizing ribose from formaldehyde (CH2O) breslow1959mechanism . Our simulations have generated a few relevant hydrocarbon products (shown in Figures S5 and S6), including formaldehyde molecule (number 1 in Figure S5), which was later chosen as the primary reactant. Additionally, to facilitate the synthesis process, the reactants also included a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5). Initially, we constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Figure S7). After $\sim$76 ps, a “C5” molecule was formed at $\sim$10 GPa and 1400 K. However, it did not transform into the structure of ribose after an extended time. We further tried to add other reactants to the simulation box, and found that after adding 5 water molecules, the ribose molecule appeared, indicating the importance role of water in the formation of ribose. The reaction process is shown in Fig. 5 (a), where the formaldehyde and “C2” molecules react to form a “C3” molecule, which then combines with another “C2” molecule to form a “C5” molecule. This “C5” molecule then evolves into the open-chain form of ribose at $\sim$19 ps.
We calculated the free energy landscape in Fig. 5 (b), where two C-C distances were selected as collective variables labelled as CV1 and CV2. For CV1, two carbon atoms come from formaldehyde and a “C2” molecule, respectively, which are labeled as C1. While for CV2, two carbon atoms are from two different “C2” molecules, marked as C2. The free energy landscape shows a clear energy minimum corresponding to the formation of ribose, but the C-C distance at this minimum is larger than a typical C-C bond distance. We examined the biased molecular dynamics trajectories and found that the carbon backbone of the generated ribose molecule has the C1-C2-C1-C2 structure. By measuring the distances of C1-C1 and C2-C2 , we confirmed that they are consistent to the distances observed in the free energy landscape. We also randomly selected another initial configuration to confirm that the ribose molecule is thermodynamically stable, as shown in Figure S8.
@@
We then shifted our attention to the chemical evolution of nucleobases, which are another essential building blocks of genetic information. There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U). Of these, uracil is unique to RNA, so our studies primarily focused on investigating the formation of uracil molecules under extreme conditions. We first studied the formation of the open-chain “C4N2” molecule which has the same CN-backbone as the open-chain form of uracil, as shown in Fig. 6 (a). Previous experimental studies on the origin of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant. However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Figure S9). Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule. The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our first-stage AIMD simulations. A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S5) and the urea molecule (CH4N2O, number 11 in Fig. 1) as reactants. The simulation results are presented in Figure S10. However, regardless of the inclusion of water molecules or the application of different pressure conditions, we did not observe any “C4N2” molecules with the CN-backbone of the open-chain form of uracil. Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figures S11, S12 and S14).
@@
Later we chose allophanic acid (C2H4N2O3, number 13 in Fig. 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c). At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule has the same CN-backbone as the open-chain form of thymine (T). As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed, which has the same CN-backbone as the open-chain form of uracil. The formation process of these two kinds of molecules are shown in Figure S15. We further performed an enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K. The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig. 6 (c). Finally, a uracil-like molecule, called uracil glycol, was synthesized, which has a uracil ring. The free energy curve in Fig. 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that the uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions. The uracil-like molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.

### v4

# Synthesis and stability of biomolecules in C-H-O-N fluids under Earth’s upper mantle conditions

@@ Abstract
How life started on Earth is an unsolved mystery. There are various hypotheses for the location ranging from outer space to the seafloor, subseafloor or potentially deeper. Here, we applied extensive ab initio molecular dynamics (AIMD) simulations to study chemical reactions between NH3, H2O, H2, and CO at pressures (P) and temperatures (T) approximating the conditions of Earth’s upper mantle (i.e. 10 - 13 GPa, 1000 - 1400 K). Contrary to the previous assumptions that large organic molecules might readily disintegrate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed without any catalysts and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, urea, and uracil-like molecules. Particularly, our free energy calculations showed that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Moreover, while the pyranose (six-membered-ring) form of ribose is more stable than the furanose (five-membered-ring) form at ambient conditions, we found that the formation of the five-membered-ring form of ribose is thermodynamically more favored at extreme conditions, which is consistent with the exclusive incorporation of $\beta$-D-ribofuranose in RNA. We have uncovered a previously unexplored pathway through which the crucial biomolecules could be abiotically synthesized from geofluids in the deep interior of Earth and other planets and these formed biomolecules could potentially contribute to the early stage of the emergency of life.
@@ I Introduction
In those hypotheses, prebiotic chemistry typically happens at non-ambient conditions. The pressure (P) of hydrothermal vents, resulting from the interaction between seafloor and seawater producing hot fluids, is tens of megapascals depending on the depth of water, and the temperature (T) can reach 700 K. Those deep sea vents were considered to have suitable conditions for the origin of life martin2008hydrothermal , but previous studies also suggested that the high temperatures in the vents could readily degrade crucial biomolecules in aqueous solutions miller1988submarine . Shock recovery experiments successfully synthesized various biomolecules, where the pressure can reach tens of gigapascals and the temperature can rise to thousands of kelvins before releasing back to ambient conditions Sekine2024 . However, due to the dynamic nature of shock waves, samples do not reach thermodynamic equilibrium states, so the shock experiments are usually conducted to mimic the conditions of meteorite impacts on Earth furukawa2009biomolecule . It was commonly assumed that large biomolecules could not persist at both high hydrostatic pressures and high temperatures as found in Earth’s upper mantle, where pressures can reach up to $\sim$13 GPa, accompanied by high temperatures of $\sim$1700 K. In recent years, Otake et al. studied the stability of amino acids and their oligomerization at 1.0–5.5 GPa and 453–673 K, and found that high-pressure conditions inhibited the decomposition of amino acids and small amino acids could be oligomerized up to pentamers. However, their experiments were conducted at the water-poor condition otake2011stability . In aqueous solutions, there seems a strong thermodynamic drive towards the formation of amino acids dimers above 1.5 GPa and temperature above 423 K robinson2021quantifying , in agreement with some experimental observations pedreira2019spontaneous . Previously, the fluids in Earth’s mantle were commonly modeled as simple mixtures of small unreactive volatile molecules, including H2O, CO2, CO, CH4, and H2 zhang2009model . However, recent experimental and theoretical studies have revealed the significant roles played by chemical speciation, aqueous ions, and complexes in the supercritical geofluids within Earth’s lithosphere manning2013chemistry ; pan2013dielectric ; sverjensky2020changing ; pan2016fate ; abramson2017water ; stolte2019large ; stolte2021water . These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules in Earth’s or other planets’ subsurface or even deeper.
@@
It is very challenging to study the building blocks of biomolecules for life in laboratory settings due to the extreme environments and subsequent analyses. Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular-level insights perez2020prebiotic . Saitta et al. applied ab initio molecular dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller . Goldman et al. applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic . Note that the shock compression simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions Ross2013comment ; Goldman2013reply . AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide . Despite the numerous previous studies, the exploration of prebiotic chemistry at both high hydrostatic pressures and high temperatures largely remains an uncharted territory.
@@
Here, we applied extensive AIMD simulations to study the reactions of inorganic C-H-O-N geofluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found in the transition zone of Earth’s upper mantle. We discovered that a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including formamide, glycine, ribose, urea, and uracil-like molecules. The formation pathways are also detailed discussed. Our free energy calculations using AIMD simulations suggest that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Our study suggests a new potential avenue for the abiotic formation of biomolecules in the deep interior of Earth and other planets, further expanding the applicable places of “primordial soup” theory.
@@ II Results and Discussion
The simulation box initially contained 15 molecules of each of the following species: H2O, H2, CO, and NH3 (Fig. S1). Those molecules are typical volatile compounds found in C-H-O-N fluids in reducing environments of Earth’s deep interior zhang2009model ; manning2013chemistry , and are also abundant in interstellar space guelin2022organic . We conducted AIMD simulations at three P-T conditions, 10 GPa and 1000 K, 10 GPa and 1400 K, and 13 GPa and 1400 K, to evaluate how chemical products vary with P and T. Those P-T conditions are typically found at the bottom of Earth’s upper mantle. After more than 400 ps simulations at each P-T condition, we found about 100 different organic species. Here, we mainly focus on the formation of biotically relevant molecules, so we classify the organic molecules into six categories: “CN”, “C2N”, “CN2”, “C2N2”, “C3Nx”, and “C4Nx”, based on the numbers of carbon and nitrogen atoms, as shown in Table 1. Fig. 1 shows the chemical structures of the representative species. Species with more than five carbon atoms were observed, however they are excluded at this stage due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds. The percents of these CN-containing species as a function of time at three different P-T conditions are shown in Fig. S3, which suggest that the concentrations of these species gradually stabilize over time.
@@
Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; saladino2015meteorite ; rios2015impact ; ferus2015high ; enchev2021chemical . Thus, we first investigated the formation of formamide molecule. Fig. 3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Fig. S4. At 10 GPa and 1000 K, the formation reaction begins with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution. The encounter between an H+ ion and a CO molecule leads to the formation of a C-H bond and an aldehyde group (-CH=O). Subsequently, the aldehyde group reacts with a NH3 molecule to generate a C-N bond, and then, one N-H bond breaks, ultimately producing formamide. At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K. One N-H bond in NH3 first breaks, leading to a dissociative amino group (-NH2). The amino group then reacts with a CO molecule, forming the C-N bond. Later, a water molecule dissociates, providing a proton that ultimately contributes to the formation of the formamide molecule. At 13 GPa and 1400 K, the C-N bond is initially formed through the reaction between NH3 and CO molecules. After a very short time, a dissociated proton from a water molecule is also bonded to the carbon atom, forming a C-H bond. Finally, by releasing a proton from the NH3 group, the formamide molecule is generated. Despite the different formation pathways at three P-T conditions, we observed two common features: first, the reactants always have CO, H2O, and NH3, and second, the reactions must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
@@
For the generation of biomolecules in living matter, the formation of C-N bonds is the key step. We further applied the advanced sampling method (see details in the Computational Methodology section) to calculate the free energy profile of the C-N bond formation at three P-T conditions in Fig. 3(b). The collective variable (CV) is the distance between the carbon and nitrogen atoms. Our calculations show that the C-N bonds at 10 GPa, 1000 K and 13 GPa, 1400 K, whose lengths are 1.33 Å and 1.31 Å  respectively, are at the metastable states. The free energy decreases upon dissociation of the C-N bond, indicating that breaking C-N bonds at these two P-T conditions are more thermodynamically favored. The free energy profiles suggest that the energy barriers of the C-N bond dissociation are 17.57 kcal/mol and 16.40 kcal/mol, at 10 GPa, 1000 K and 13 GPa, 1400 K, respectively. As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 49.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition. The free energy result agrees with the data in Fig. 2, where there are more CN-containing species at 10 GPa and 1400 K than at the other two P-T conditions. Additionally, compared with the free energy profile at 10 GPa and 1000 K, there are two more energy minima at the C-N distance of 2.37 Å  and 2.45 Å at 10 GPa, 1400 K and 13 GPa, 1400 K, respectively. The free energy increases upon the dissociation of the C-N pair, indicating that these two minima are thermodynamically stable. The C-N distances are larger than that of a typical C-N covalent bond, which implies the presence of an additional atom connecting the two. This is consistent with a wider variety of organic species at these two P-T conditions. Overall, our free energy calculations suggest that the C-N bond formation is favored at 10 GPa and 1400 K.
@@
After studying the basic chemical products at three P-T conditions, we now turn to investigate the formation of large biomolecules as potential building blocks of life. Previous simulations have extensively studied glycine and the smallest amino acid under various conditions saitta2014miller ; goldman2010synthesis ; wang2014discovering . In our simulations at 10 GPa and 1400 K, we discovered $\alpha$-hydroxyglycine, very similar to glycine. Fig. 4 (a) illustrates the process of forming $\alpha$-hydroxyglycine from formamide. A proton attacks the formamide molecule, leading to the creation of a protonated formamide molecule. Then, the protonated formamide reacts with a large CO2-containing molecule, resulting in the “-N-C-C-” backbone through forming a C-C bond. Finally, another proton is captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule. In comparison to glycine, $\alpha$-hydroxyglycine differs by having one hydrogen atom and an “-OH” group instead of two H atoms connecting to $\alpha$-C. Glycine can be obtained if the “-OH” group is replaced by a hydrogen atom. During this process, the formation of the C-H bond is a crucial step. To explore appropriate P-T conditions for the possible synthesis of glycine, we calculated the free energy as a function of the C-H distance in Fig. 4 (b). Considering that the CN-backbone of the “C2N” species at 13 GPa - 1400 K is “-C-N-C-”, different from the CN-backbone in the glycine molecule, we did not consider this P-T condition. We can see that the free energy minima is at 1.07 Å for both 10 GPa, 1000 K and 10 GPa, 1400 K, and the energy barriers to break the C-H bond are 21.91 kcal/mol and 22.14 kcal/mol, respectively, which are very similar. After the C-H bond dissociation, the free energy increases more at 10 GPa and 1400 K (6.76 kcal/mol) than at 10 GPa and 1000 K (2.63 kcal/mol), indicating that the C-H bond is more favored at the former P-T condition.
@@
To study the reaction from $\alpha$-hydroxyglycine to glycine at extreme P-T conditions, we set up a new simulation box containing 5 $\alpha$-hydroxyglycine, 10 H2, and 10 H2O molecules, as shown in Fig. 4(c) and performed AIMD simulations at 10 GPa and 1400 K. Initially, two $\alpha$-hydroxyglycine molecules combine to form one molecule, but later dissociate to dehydroglycine and glyoxylic acid after $\sim$28 ps. As the simulation goes on, the dehydroglycine will be reduced by H2 and finally it becomes glycine at $\sim$ 38 ps.
@@
We first considered ribose (C5H10O5). The formose reaction is a well-known prebiotic pathway for synthesizing ribose from formaldehyde (CH2O) breslow1959mechanism . Our simulations have generated a few relevant hydrocarbon products (shown in Figs. S5 and S6), including formaldehyde molecule (number 1 in Fig. S5), which was later chosen as the primary reactant. Additionally, to facilitate the synthesis process, the reactants also included a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Fig. S5). Initially, we constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Fig. S7). After $\sim$76 ps, a “C5” molecule was formed at $\sim$10 GPa and 1400 K. However, it did not transform into the structure of ribose after an extended time. We further tried to add other reactants to the simulation box, and found that after adding 5 water molecules, the ribose molecule appeared, indicating the importance role of water in the formation of ribose. The reaction process is shown in Fig. 5 (a), where the formaldehyde and “C2” molecules react to form a “C3” molecule, which then combines with another “C2” molecule to form a “C5” molecule. This “C5” molecule then evolves into the open-chain form of ribose at $\sim$19 ps.
We calculated the free energy landscape in Fig. 5 (b), where two C-C distances were selected as collective variables labelled as CV1 and CV2. For CV1, two carbon atoms come from formaldehyde and a “C2” molecule, respectively, which are labeled as C1. While for CV2, two carbon atoms are from two different “C2” molecules, marked as C2. The free energy landscape shows a clear energy minimum corresponding to the formation of ribose, but the C-C distance at this minimum is larger than a typical C-C bond distance. We examined the biased molecular dynamics trajectories and found that the carbon backbone of the generated ribose molecule has the C1-C2-C1-C2 structure. By measuring the distances of C1-C1 and C2-C2 , we confirmed that they are consistent to the distances observed in the free energy landscape. Furthermore, another initial configuration for the free energy calculation is also presented in Fig. S8, the obtained free energy landscape still shows that the ribose molecule is thermodynamically stable.
@@
We then shifted our attention to the chemical evolution of nucleobases, which are another essential building blocks of genetic information. There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U). Of these, uracil is unique to RNA, so our studies primarily focused on investigating the formation of uracil molecules under extreme conditions. We first studied the formation of the open-chain “C4N2” molecule which has the same CN-backbone as the open-chain form of uracil, as shown in Fig. 6 (a). Previous experimental studies on the origin of life suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant. However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Fig. S9). Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule. The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our first-stage AIMD simulations. A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Fig. S5) and the urea molecule (CH4N2O, number 11 in Fig. 1) as reactants. The simulation results are presented in Fig. S10. However, regardless of the inclusion of water molecules or the application of different pressure conditions, we did not observe any “C4N2” molecules with the CN-backbone of the open-chain form of uracil. Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figs. S11, S12 and S14).
@@
Later we chose allophanic acid (C2H4N2O3, number 13 in Fig. 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Fig. S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c). At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps. This molecule has the same CN-backbone as the open-chain form of thymine (T). As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed, which has the same CN-backbone as the open-chain form of uracil. We further performed an enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K. The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig. 6 (c). Finally, a uracil-like molecule, called uracil glycol, was synthesized, which has a uracil ring. The free energy curve in Fig. 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that the uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions. The formation process and reaction mechanism of the uracil-like molecule are shown in Fig. S15. The uracil-like molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.